OpenStudy (anonymous):

What is the perimeter of the triangle formed by the x-axis, y-axis, and the line 4x+12=3y? 7 9 12 13 16

OpenStudy (ghazi):

find the x and y intercept of this line..put x=0 for y and y=0 for x

OpenStudy (anonymous):

x=0 for y and y=0 for x?

OpenStudy (ghazi):

to find the x and y intercept put x=0 to find y intercept and to find x intercept put y=0 you'll have x=-3 and y=4

OpenStudy (anonymous):

so which choice would it be?

OpenStudy (ghazi):

now find another side of triangle....it will be hypotenuse =\[\sqrt{(-3)^2+(4)^2}\]=5 and perimeter = 5+3+4=12

OpenStudy (anonymous):

thnx!

OpenStudy (ghazi):

:)